How many gallons of 65% antifreeze and how many gallons of 20% antifreeze should be mixed to obtain 50 gallons of a 56% mixture of antifreeze? (Hint: Let x represent the number of gallons of 65% antifreeze. Then 50 - x represents the amount of 20% antifreeze.)

Answer by TimothyLamb(4379)

65s + 20w = 56(50)
s + w = 50
w = 50 - s
---
65s + 20w = 56(50)
65s + 20(50 - s) = 2800
65s + 1000 - 20s = 2800
45s = 1800
---
s = 40 gal
w = 10 gal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php